OpenStudy (anonymous):

how do i simplify (-2+2/3x)6

OpenStudy (anonymous):

if I am reading correctly it would reduce to (0/3x)6 which is 0 * 6 = 0

jhonyy9 (jhonyy9):

((-6x+2)/3x)6=(-6x+2)6/3x=(-6x+2)2/x=(-12x+4)/x

jhonyy9 (jhonyy9):

sorry if x not is denominator --- so will be (-6+2)/3x)6=(-4/3x)6=(-4x)6/3=(-4x)2= -8x

OpenStudy (anonymous):

thanks for the help, have a problem with dealing with fractions